A bullet going with speed 350 m/s enters a concrete wall and penetrates a distance of 5.0 cm before coming to rest. Find the deceleration.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. . u = 350 m/s, s = 5 cm = 0.05 m, v = 0 Deceleration = a = (V^2-u^2)/2s = (0-(350)^2)/(2 x 0.05) = -12.22 x 105 m/s2. Deceleration is 12.2 × 105 m/s2.